. A 25mm diameter rod is subjected to an axial load of 20 000N. If the material is elastic with a Young's Modulus of Elasticity E = 70,000Nmm-2 estimate the percentage elongation.
1
Expert's answer
2020-06-10T18:23:05-0400

Solution.

"d=25mm;"

"F=20000N;"

"E=70000N\/mm^2;"

"\sigma=\dfrac{F}{S};"

"S=\dfrac{\pi d^2}{4};"

"\sigma=E\epsilon\implies\epsilon=\dfrac{\sigma}{E};"


"S=3.14\sdot\dfrac{(25mm)^2}{4}=490.625mm^2;"

"\sigma=\dfrac{20000N}{490.625mm^2}=40.76N/mm^2;"

"\epsilon=\dfrac{40.76N/mm^2}{70000N/mm^2}=0.00058=0.058\%;"

Answer: "\epsilon=0.058\%."